# Connectedness of fibers beyond semitoric systems I: the non-degenerate case

### v3

@@ Abstract.
In this paper we study the connectedness of the fibers of integrable systems that have only non-degenerate singular points and extend complexity one $T$-spaces with proper moment maps. Our main result states that if there are no tall singular points with a hyperbolic block and connected $T$-stabilizer, then each fiber is connected. Moreover, we prove that the above condition is necessary if either some reduced space is simply connected or the moment map for the integrable system is generic in a natural sense.
@@ 1. Introduction
An important step in studying Hamiltonian group actions is understanding the fibers of the moment map. In this direction, connectedness of the fibers has been established in the following cases:
@@
In both cases, the connectedness of the fibers of the moment map is key to understanding the moment map image and plays an important role in classification.
In this paper we study the connectedness of the fibers of integrable systems that are extensions of complexity one torus actions, and have only non-degenerate singular points. Complexity one torus actions are the simplest after the toric case and have been partially classified in [KT1, KT2, KT3]. Non-degenerate singular points of integrable systems are generic, and can be thought of as the symplectic analog of critical points for Morse-Bott functions; they have attracted much interest, both for classical and quantum integrable systems (see [BF, SV] and references therein). In particular, the integrable systems that we study include the semi-toric systems of [VN], and their higher dimensional generalization in [Wa].
In our main result, Theorem 1.6 we give sufficient conditions for fibers of the moment map to be connected, generalizing the results in [VN, Wa]. Morse theory plays a central role in the theorems (1) and (2) above, and our theorem both uses and generalizes the following simple Morse-theoretic result: Let $N$ be a closed, connected 2-dimensional manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a Morse function. If $\overline{g}$ has no singular point of index one, then each fiber is connected (see Lemma 7.3). We also explore the extent to which the hypotheses in our main result are necessary (see Theorem 1.10).
@@
Let $(M,\omega,\Phi)$ be a complexity one $T$-space. Given $\beta\in{\mathfrak{t}}^{*}$, the reduced space at $\beta$ is the quotient $\Phi^{-1}(\beta)/T$; if $\beta=0$, we write M//T{M/\!/T} for the reduced space. A point $p\in M$ is short if $[p]$ is an isolated point in the reduced space $\Phi^{-1}(\Phi(p))/T$ and is tall otherwise. By [KT1, Proposition 6.1], if $\Phi$ is proper each reduced space $\Phi^{-1}(\beta)/T$ containing more than one point is topologically a closed, connected oriented 2-dimensional manifold; moreover, any two non-empty reduced spaces have the same fundamental group [KT1, Lemma 5.7 and Corollary 9.7].
The main objects that we study in this paper are non-degenerate integrable systems that extend complexity one torus actions, i.e., integrable systems $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ such that every singular point is non-degenerate and such that $(M,\omega,\Phi)$ is a complexity one $T$-space. Our main result is as follows.
@@
  • The reduced space $\Phi^{-1}(\beta_{0})/T$ is simply connected for some $\beta_{0}\in\Phi(M).$
@@
Then the fiber $f^{-1}(\beta,c)$ is connected for all $(\beta,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$ if and only $M$ has no tall singular points with a hyperbolic block and connected $T$-stabilizer.
@@
If in Theorems 1.6 and 1.10 we additionally assume that $M$ is compact, then by a theorem of Hui Li [Li], the fundamental group of $M$ is equal to that of any reduced space. Hence, in this case, the phrase “the reduced space $\Phi^{-1}(\beta)/T$” can be changed to “MM”.
The fundamental idea behind the proofs of Theorems 1.6 and 1.10 is quite simple: Since $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g)\right)$ is an integrable system, the function $g$ is $T$-invariant, and hence induces a function $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$. By the key technical result in our paper, Proposition 7.1, each non-trivial component of the reduced space $\Phi^{-1}(\beta)/T$ can be given the structure of a smooth oriented surface so that $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$ is a Morse function. Moreover, the critical points of index $1$ correspond exactly to orbits of non-degenerate singular points with a hyperbolic black and connected $T$-stabilizer. Since $\Phi$ is proper, the reduced space $\Phi^{-1}(\beta)/T$ is compact and connected by Theorem 5.5. Therefore, the proofs reduce to standard results in Morse theory; see Lemmas 7.3 and 7.4.
This method can be used more generally on integrable systems that extend a complexity one torus action but have (sufficiently nice) degenerate singular points. For example, in a forthcoming paper, we will prove that the analogs of Theorems 1.6 and 1.10 hold if we also allow a broader class of “ephemeral” singular points; like focus-focus points and hyperbolic points with disconnected $T$-stabilizers, these correspond to regular points on the symplectic quotients.
@@
More generally, Theorems 1.6 and 1.10 both hold if $\Phi$ is proper as a map to an open convex subset of $\mathfrak{t}$$^{*}$, because the analog of Theorem 5.5 still holds in this case [LMTW, Theorem 4.3]. Additionally, if $f$ is proper and the fibers of $\Phi$ are connected, then the first statement in Theorem 1.6 still holds. This is because a proper Morse function on a connected surface with no critical points of index $1$ has connected fibers (Proposition 5.6).
Remark 1.14.
Let $(M,\omega,{\mathfrak{t}}^{*}\times{\mathbb{R}},f)$ be an integrable system satisfying the assumptions in the first two sentences of Theorem 1.6, but assume that the reduced space $\Phi^{-1}(\beta)/T$ has positive genus $k$, and hence Euler characteristic 22k2-2k. In this case, by calculating the Euler characteristic, we see that $\Phi^{-1}(\beta)/T$ has at least $2k$ orbits of non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer because because $\overline{g}$ has at least one minimum and maximum. Note that if $\Phi^{-1}(\beta)/T$ has more than $2k$ such orbits, then the same argument shows that $\overline{g}$ must have additional local minima or maxima; hence the fibers of $f$ are not connected by Lemma 7.2.
Structure of the paper
In Section 2, we review some facts about tall complexity one local models and then briefly explore a few additional properties. Next, in Section 3 we prove that we can endow the symplectic quotient $\Phi^{-1}(\beta)/T$ with a smooth structure so that $\overline{g}$ is a Morse function – as long as most singular points satisfy a certain technical condition that guarantees that we can do this locally. Most of the rest of the paper is dedicated to proving that non-degenerate singular points satisfy this condition. For example, in Section 4 we prove that tall points that have purely elliptic type satisfy it, and in Section 5 we show that, if $M$ is four-dimensional, focus-focus points satisfy it. At this point, we are able to demonstrate our technique by using it to give a new proof that that fibers of a semi-toric system are connected. In Section 6, we prove that all the remaining types of non-degenerate singular points also satisfy this technical condition. Therefore, we can complete the proof of our theorems in Section 7.
Acknowledgments
D. Sepe was partially supported by FAPERJ grant JCNE E-26/202.913/2019 and by a CAPES/Alexander von Humboldt Fellowship for Experienced Researchers 88881.512955/2020-01. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brazil (CAPES) – Finance code 001. S. Tolman was partial supported by NSF DMS Award 2204359 and Simons Foundation Collaboration Grant 637995.
2. Tall local models
To begin, we review some facts about tall complexity one local models. In particular, we introduce (tall) complexity one local models, construct the “defining polynomial” for tall local models, and use it to describe the symplectic quotient Y//T{Y/\!/T}. Moreover, we provide a generating set for $T$-invariant polynomials on these models.
Set $Y:=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, where $\mathfrak{h}$$^{\circ}$ is the annihilator of $\mathfrak{h}$} in $\mathfrak{t}$$^{*}$ and $H$ acts on $T$ by multiplication, on $\mathfrak{h}$$^{\circ}$ trivially, and on ${\mathbb{C}}^{h+1}$ by $h\cdot z=\rho(h^{-1})z$. Fix an inner product on the Lie algebra $\mathfrak{t}$} once and for all, and use it to identify $\mathfrak{t}$$^{*}$ with ${\mathfrak{h}}^{\circ}\oplus{\mathfrak{h}}^{*}.$ There exists a canonical symplectic form $\omega_{Y}$ on $Y$ so that the $T$-action $s\cdot[t,\alpha,z]=[st,\alpha,z]$ has homogeneous moment map $\Phi_{Y}\colon Y\to{\mathfrak{t}}^{*}$ given by
$\Phi_{Y}([t,\alpha,z])=\alpha+\Phi_{H}(z),$
where $\Phi_{H}(z)=\frac{1}{2}\sum_{i}\eta_{i}|z_{i}|^{2}$ is the homogeneous moment map for the $H$-action on ${\mathbb{C}}^{h+1}$; see Remark 4.2. By the Marle-Guillemin-Sternberg local normal form theorem, there is a $T$-equivariant symplectomorphism from an invariant neighborhood of $\mathfrak{O}$ to an open subset of $Y$ taking $p$ to $[1,0,0]$. We call $(Y,\omega_{Y},\Phi_{Y})$ the local model for $p$. We say that $Y$ is short if $\Phi_{Y}^{-1}(0)$ consists of a single $T$-orbit and tall otherwise. By Lemma 2.7 below, the fiber $\Phi^{-1}_{Y}(0)$ contains exactly one $T$-orbit exactly if $p]\in\Phi^{-1}(\Phi(p))/T$ is an isolated point. Hence, $p$ is short exactly if its local model is short.
As shown in [KT1], tall local models [^1] Here, and throughout the paper, all the local models that we consider are complexity one. have an important invariant, called the defining polynomial.
@@ 2. Tall local models
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a local model associated to an injective homomorphism $\rho\colon H\to(S^{1})^{h+1}$. Then there exists $\xi=(\xi_{0},\ldots,\xi_{h})\in{\mathbb{Z}}^{h+1}$ such that the homomorphism $(S^{1})^{h+1}\to S^{1}$ given by $\lambda\mapsto\prod\limits_{j=0}^{h}\lambda_{j}^{\xi_{j}}$ induces a short exact sequence
@@
Existence of $\xi$ and its uniqueness up to sign follow from [KT1, Lemma 5.8]. The local model $Y$ is tall if and only if $\Phi^{-1}_{Y}(0)$ consists of more than one orbit. By [KT1, Lemma 5.4], the latter occurs exactly if $\Phi_{Y}$ is not proper, which, by [KT1, Lemma 5.8], is equivalent to $\xi_{i}\xi_{j}\geq 0$ for all $i,j$. ∎
@@
Fix a tall local model $Y$. By Lemma 2.1, there exists a unique $\xi=(\xi_{0},\ldots,\xi_{h})\in{\mathbb{Z}}_{\geq 0}^{h+1}$ inducing the short exact sequence (2.2). The map $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$ given by $P(z)=\prod_{j=0}^{h}z_{j}^{\xi_{j}}$ is called the defining polynomial of $Y$. By a slight abuse of notation, we use the same name for the map $P\colon Y\to{\mathbb{C}}$ given by $t,\alpha,z]\mapsto\prod_{j=0}^{h}z_{j}^{\xi_{j}}$. The sum $N:=\sum_{j=0}^{h}\xi_{j}$ is called the degree of $P$.
@@
Let ${\mathbb{Z}}_{2}=\{\pm 1\}$ act on $\mathbb{C}$} by $\lambda\cdot z=\lambda z$. The associated local model $S^{1}\times_{{\mathbb{Z}}_{2}}{\mathbb{R}}^{*}\times{\mathbb{C}}$ is tall. The defining polynomial $P\colon{\mathbb{C}}\to{\mathbb{C}}$ is given by $P(z)=z^{2}$ and has degree $2$.
@@
Let $S^{1}$ act effectively on ${\mathbb{C}}^{2}$ by $\lambda\cdot(z_{1},z_{2})=(\lambda z_{1},\lambda^{-1}z_{2})$ with moment map $\Phi(z_{1},z_{2})=\frac{1}{2}(|z_{1}|^{2}-|z_{2}|^{2})$. This is a tall local model. The defining polynomial $P\colon{\mathbb{C}}^{2}\to{\mathbb{C}}$ is given by $P(z_{1},z_{2})=z_{1}z_{2}$ and also has degree $2$.
@@
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with moment map $\Phi_{Y}\colon Y\to{\mathfrak{t}}^{*}$. The defining polynomial $P\colon Y\to{\mathbb{C}}$ induces a homeomorphism $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ that restricts to a diffeomorphism on the complement of $\left\{[1,0,0]\right\}$.
@@
Following [KT1], we say that an orbit is exceptional if every nearby orbit in the same moment fiber has strictly smaller stabilizer. Arguing as in the proof of Lemma 2.1, since $Y$ is tall the moment map $\Phi_{Y}$ is not proper. The first claim follows at once by [KT1, Corollary 6.3]. By [KT1, Lemmas 5.15 and 5.17], the orbit of $[1,0,0]$ is the only possible exceptional orbit in Y//T{Y/\!/T}. The second claim now follows at once by [KT1, Corollary 7.2]. ∎
In most cases, the homeomorphism $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is not globally a diffeomorphism; however, as the lemma below shows, it is a diffeomorphism when $P(z)=z_{0}$.
@@
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with defining polynomial $P(z)=\prod_{i=0}^{h}z_{i}^{\xi_{i}}$. Assume that $\xi_{i}=0$ for all $i>k$. Let the group
@@
induces a diffeomorphism between ${\mathbb{C}}^{k+1}/\!/K$ and Y//T{Y/\!/T}. In particular, if $P(z)=z_{0}$ then $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is a diffeomorphism with inverse $z\mapsto[1,0,(z,0)].$

### v4

@@ Abstract.
In this paper we study the connectedness of the fibers of integrable systems that extend complexity one $T$-spaces with proper moment maps, assuming that every tall singular point is non-degenerate. Our main result states that if there are no tall singular points with a hyperbolic block and connected $T$-stabilizer, then each fiber is connected. Moreover, we prove that the above condition is necessary if either some reduced space is simply connected or the moment map for the integrable system is generic in a natural sense.
@@ 1. Introduction
A crucial step in studying Hamiltonian group actions is understanding the fibers of the moment map. In this direction, connectedness of the fibers has been established in the following cases:
@@
In both cases the connectedness of the fibers of the moment map is the first step towards classification results: For instance, it is central to proving convexity of the moment map image of Hamiltonian torus actions (see [At, GS2, LMTW]), and to obtaining a convexity result for semi-toric systems in dimension four (see [VN, Wa]).
In this paper we study the connectedness of the fibers of integrable systems that are extensions of complexity one torus actions, assuming that every tall singular point is non-degenerate. We hope that this will eventually allow us to extend the classification of semi-toric systems in dimension four to higher dimensions. Complexity one torus actions are the simplest after the toric case and have been partially classified in [KT1, KT2, KT3]. Non-degenerate singular points of integrable systems are generic, and can be thought of as the symplectic analog of critical points of Morse-Bott functions; they have attracted much interest, both for classical and quantum integrable systems (see [BF, SV] and references therein). In particular, the integrable systems that we study include the semi-toric systems of [VN], and their higher dimensional generalization in [Wa].
In our main result, Theorem 1.6, we give sufficient conditions for fibers of the moment map to be connected, generalizing the results in [VN, Wa]. Morse theory plays a central role in the theorems (1) and (2) above, and our theorem both uses and generalizes the following simple Morse-theoretic result: Let $N$ be a closed, connected 2-dimensional manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a Morse function. If $\overline{g}$ has no singular point of index one, then each fiber is connected (see Lemma 7.3). We also explore the extent to which the hypotheses in our main result are necessary (see Theorem 1.11).
@@
In particular, by definition regular points are non-degenerate and have purely elliptic type.
@@
Let $(M,\omega,\Phi)$ be a complexity one $T$-space. Given $\beta\in{\mathfrak{t}}^{*}$, the reduced space at $\beta$ is the quotient $\Phi^{-1}(\beta)/T$, endowed with the subquotient topology; if $\beta=0$, we write M//T{M/\!/T} for the reduced space. A point $p\in M$ is short if $[p]$ is an isolated point in the reduced space $\Phi^{-1}(\Phi(p))/T$ and is tall otherwise. By [KT1, Proposition 6.1], if $\Phi$ is proper each reduced space $\Phi^{-1}(\beta)/T$ containing more than one point is homeomorphic to (but not necessarily diffeomorphic to) a closed, connected oriented 2-dimensional manifold; moreover, any two non-empty reduced spaces have the same fundamental group [KT1, Lemma 5.7 and Corollary 9.7].
The main objects that we study in this paper are integrable systems that extend complexity one torus actions and have the property that every tall singular point is non-degenerate. Our main result is as follows.
@@
In addition to semi-toric systems, Theorem 1.6 applies to the following well-known integrable system.
Example 1.7.
The 1:21\!\!:\!\!2 resonance is the integrable system $({\mathbb{C}}^{2},\omega,{\mathbb{R}}\times{\mathbb{R}},f:=(\Phi,g))$, where $\omega\in\Omega^{2}({\mathbb{C}}^{2})$ is the standard symplectic form,
$\Phi(z_{1},z_{2}):=\textstyle\frac{1}{2}|z_{1}|^{2}+|z_{2}|^{2},\quad\text{and}\quad g(z_{1},z_{2}):={\mathfrak{R}}(z^{2}_{1}\overline{z}_{2}).$
As explained in [CVN, Sections 3.3.2 and 3.3.3], the origin, which is the only short point, is singular and degenerate. Every other singular point in this integrable system is non-degenerate; moreover, $z\in{\mathbb{C}}^{2}$ is a singular point with a hyperbolic block exactly if $z\in\{0\}\times{\mathbb{C}}^{\times}$, that is, exactly if its $S^{1}$-stabilizer is ${\mathbb{Z}}_{2}=\{\pm 1\}$. Therefore, the 1:21\!\!:\!\!2 resonance satisfies the hypothesis of Theorem 1.6. In particular, the fibers of $f$ are connected and each reduced space of the $S^{1}$-action is simply connected (see [CVN, Remark 3.1]).
@@
  • The reduced space $\Phi^{-1}(\beta_{0})/T$ is simply connected for some (and hence all) $\beta_{0}\in\Phi(M).$
@@
Then the fiber $f^{-1}(\beta,c)$ is connected for all $(\beta,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$ if and only if no tall singular point in $M$ has a hyperbolic block and connected $T$-stabilizer.
@@
If $M$ is simply connected, then the first bullet point in Theorem 1.11 automatically holds. To see this, first note that the quotient space $M/T$ is simply connected by, e.g., [B, Chapter II, Corollary 6.3]. The claim then follows by [Li2, Theorem 1.5]. If $M$ has fixed points then, conversely, under the assumptions of Theorem 1.6 the manifold $M$ is simply connected by [Li2, Theorems 1.4 and 1.5].
Remark 1.13.
More generally, Theorems 1.6 and 1.11 both hold if we replace $\Phi$ is proper” by $\Phi$ is proper as a map to an open convex subset of $\mathfrak{t}$$^{*}$”. To see this, simply replace Theorem 5.5 in the proofs by its generalization [LMTW, Theorem 4.3]. Similarly, if we replace $\Phi$ is proper” by “$f$ is proper and the fibers of $\Phi$ are connected”, then the first claim in Theorem 1.6 still holds; see Proposition 7.5.
The fundamental idea behind the proofs of Theorems 1.6 and 1.11 is quite simple: Since $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g)\right)$ is an integrable system, the function $g$ is $T$-invariant, and hence induces a function $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$. The key technical result in our paper is that we can choose a smooth structure on the reduced space $\Phi^{-1}(\beta)/T$ so that each non-trivial component is a smooth oriented surface and $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$ is a Morse function; moreover, the critical points of index $1$ correspond exactly to orbits of tall singular points with a hyperbolic block and connected $T$-stabilizer. This is stated in Proposition 7.1, which combines Propositions 3.6, 4.4 and 6.3. Since $\Phi$ is proper, the reduced space $\Phi^{-1}(\beta)/T$ is compact and connected [At, GS2, LMTW]. Therefore, the proofs reduce to standard results in Morse theory; see Lemmas 7.3 and 7.4.
This method can be used more generally on integrable systems that extend a complexity one torus action but have (sufficiently nice) degenerate singular points. For example, in a forthcoming paper [ST], we prove that the analogs of Theorems 1.6 and 1.11 hold if we also allow systems with a broader class of “ephemeral” singular points, including more general resonances than that in Example 1.7. Like focus-focus points and hyperbolic points with disconnected $T$-stabilizers, these ephemeral singular points correspond to regular points on the symplectic quotients.
Remark 1.14.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space with a proper moment map. Assume that each tall singular point is non-degenerate. However, suppose that the reduced space $\Phi^{-1}(\beta)/T$ has positive genus $k$. (For example, let $M$ be the product of a symplectic toric manifold with a closed, orientable surface of genus $k$, the map $\Phi$ be (the pullback of) the moment map for the toric action, and $g$ be (the pullback of) a Morse function on the surface.) In this case, by a variation on the above argument, $\Phi^{-1}(\beta)/T$ has at least $2k$ orbits of non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer. Additionally, if $\Phi^{-1}(\beta)/T$ has more than $2k$ such orbits, then $\overline{g}$ must have additional local minima or maxima; hence the fibers of $f$ are not connected by Morse theory; see Lemma 7.2.
@@
Structure of the paper
In Section 2, we review some facts about tall complexity one local models and then briefly explore a few additional properties. Next, in Section 3 we prove that we can endow the symplectic quotient $\Phi^{-1}(\beta)/T$ with a smooth structure so that $\overline{g}$ is a Morse function – as long as most singular points satisfy a certain technical condition that guarantees that we can do this locally. Most of the rest of the paper is dedicated to proving that non-degenerate singular points satisfy this condition. For example, in Section 4 we prove that tall points that have purely elliptic type satisfy it, and in Section 5 we show that, if $M$ is four-dimensional, focus-focus points satisfy it. At this point, we are able to demonstrate our technique by using it to give a new proof that fibers of a semi-toric system are connected. In Section 6, we prove that all the remaining types of non-degenerate singular points also satisfy this technical condition. Therefore, we can complete the proof of our theorems in Section 7.
Acknowledgments
D. Sepe was partially supported by FAPERJ grant JCNE E-26/202.913/2019 and by a CAPES/Alexander von Humboldt Fellowship for Experienced Researchers 88881.512955/2020-01. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brazil (CAPES) – Finance code 001. S. Tolman was partially supported by NSF DMS Award 2204359 and Simons Foundation Collaboration Grant 637995.
2. Tall local models
In this section we review some facts about tall complexity one local models. In particular, we introduce (tall) complexity one local models, construct the “defining monomial” for tall local models, and use it to describe the symplectic quotient Y//T{Y/\!/T}. Moreover, we provide a generating set for $T$-invariant polynomials on these models.
Set $Y:=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$, where $\mathfrak{h}$$^{\circ}$ is the annihilator of $\mathfrak{h}$} in $\mathfrak{t}$$^{*}$ and $H$ acts on $T$ by multiplication, on $\mathfrak{h}$$^{\circ}$ trivially, and on ${\mathbb{C}}^{h+1}$ by $h\cdot z=\rho(h^{-1})z$. Fix an inner product on the Lie algebra $\mathfrak{t}$} once and for all, and use it to identify $\mathfrak{t}$$^{*}$ with ${\mathfrak{h}}^{\circ}\oplus{\mathfrak{h}}^{*}.$ There exists a canonical symplectic form $\omega_{Y}$ on $Y$ so that the $T$-action $s\cdot[t,\alpha,z]=[st,\alpha,z]$ has homogeneous moment map $\Phi_{Y}\colon Y\to{\mathfrak{t}}^{*}$ given by
$\Phi_{Y}([t,\alpha,z])=\alpha+\Phi_{H}(z),$
where $\Phi_{H}(z)=\frac{1}{2}\sum_{i}|z_{i}|^{2}\eta_{i}$ is the homogeneous moment map for the $H$-action on ${\mathbb{C}}^{h+1}$; see Remark 2.2. By the Marle-Guillemin-Sternberg local normal form theorem, there is a $T$-equivariant symplectomorphism from an invariant neighborhood of $\mathfrak{O}$ to an open subset of $Y$ taking $p$ to $[1,0,0]$.
Definition 2.1.
The triple $(Y,\omega_{Y},\Phi_{Y})$ is the local model for $p$.
@@ 2. Tall local models
To construct the symplectic form on $Y$, let $i^{*}\colon{\mathfrak{t}}^{*}\to{\mathfrak{h}}^{*}$ be the projection induced by the inclusion $H\hookrightarrow T$. Equip $T^{*}T\times{\mathbb{C}}^{h+1}$ with the symplectic form given by the sum of the pullbacks of the canonical symplectic form on $T^{*}T$ and the standard symplectic form on ${\mathbb{C}}^{h+1}$. The $H$-action on $T^{*}T\times{\mathbb{C}}^{h+1}\cong T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}$ given by $h\cdot(t,\beta,w):=(ht,\beta,\rho(t^{-1})w)$ has a moment map $T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ that takes $(t,\beta,z)$ to $i^{*}(\beta)-\Phi_{H}(z)$. The reduced space at 0 is smooth, and the map $Y\to(T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1})/\!/H$ taking $t,\alpha,z]$ to $t,\alpha+\Phi_{H}(z),z]$ is a diffeomorphism. By definition, the symplectic form $\omega_{Y}$ on $Y$ is the pullback of the reduced symplectic form along this diffeomorphism.
We say that the local model $Y$ is short (respectively tall) if $1,0,0]\in Y$ is short (respectively tall). As shown in [KT1], tall local models [^1] Here, and throughout the paper, all the local models that we consider are complexity one. have an important invariant, called the defining monomial.
@@
Let $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$ be a local model associated to an injective homomorphism $\rho\colon H\to(S^{1})^{h+1}$. Then there exists $\xi=(\xi_{0},\ldots,\xi_{h})\in{\mathbb{Z}}^{h+1}$ such that the homomorphism $(S^{1})^{h+1}\to S^{1}$ given by $\lambda\mapsto\prod\limits_{j=0}^{h}\lambda_{j}^{\xi_{j}}$ induces a short exact sequence
@@
Existence of $\xi$ and its uniqueness up to sign follow from [KT1, Lemma 5.8]. By [KT1, Lemma 5.4] (and its proof), the local model $Y$ is tall exactly if $\Phi_{Y}$ is not proper, which, by [KT1, Lemma 5.8], is equivalent to $\xi_{i}\xi_{j}\geq 0$ for all $i,j$. ∎
@@
Fix a tall local model $Y$. By Lemma 2.3, there exists a unique $\xi=(\xi_{0},\ldots,\xi_{h})\in{\mathbb{Z}}_{\geq 0}^{h+1}$ inducing the short exact sequence (2.4). The map $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$ given by $P(z)=\prod_{j=0}^{h}z_{j}^{\xi_{j}}$ is called the defining monomial of $Y$. By a slight abuse of notation, we use the same name for the map $P\colon Y\to{\mathbb{C}}$ given by $t,\alpha,z]\mapsto\prod_{j=0}^{h}z_{j}^{\xi_{j}}$. The sum $N:=\sum_{j=0}^{h}\xi_{j}$ is called the degree of $P$.
@@
Let ${\mathbb{Z}}_{2}=\{\pm 1\}$ act on $\mathbb{C}$} by $\lambda\cdot z=\lambda z$. The associated local model $S^{1}\times_{{\mathbb{Z}}_{2}}{\mathbb{R}}^{*}\times{\mathbb{C}}$ is tall. The defining monomial $P\colon{\mathbb{C}}\to{\mathbb{C}}$ is given by $P(z)=z^{2}$ and has degree $2$.
@@
Let $S^{1}$ act effectively on ${\mathbb{C}}^{2}$ by $\lambda\cdot(z_{1},z_{2})=(\lambda z_{1},\lambda^{-1}z_{2})$ with moment map $\Phi(z_{1},z_{2})=\frac{1}{2}(|z_{1}|^{2}-|z_{2}|^{2})$. This is a tall local model. The defining monomial $P\colon{\mathbb{C}}^{2}\to{\mathbb{C}}$ is given by $P(z_{1},z_{2})=z_{1}z_{2}$ and also has degree $2$.
As in [KT1], we say that a point $p\in M$ (and the orbit $p]\in M/T$) is exceptional if every point in some neighborhood of $[p]$ in the reduced space $\Phi^{-1}(\Phi(p))/T$ has strictly smaller stabilizer. By definition, every short point is exceptional. The following criterion determines if a point in a tall local model is exceptional.
Lemma 2.9.
Fix a tall local model $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$ with defining monomial $P(z)=\prod_{i=0}^{h}z_{i}^{\xi_{i}}$. Then $t,\beta,z]\in Y$ is exceptional exactly if $\sum_{z_{i}=0}\xi_{i}>1$, where the sum is over all $i$ such that $z_{i}=0$.
Proof.
Arguing as in the proof of Lemma 2.3, since $Y$ is tall the moment map $\Phi_{Y}$ is not proper. Therefore, the claim can be obtained by joining the statements of [KT1, Lemmas 5.15 and 5.17]. ∎
@@
Let $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$ be a tall local model with moment map $\Phi_{Y}\colon Y\to{\mathfrak{t}}^{*}$. The defining monomial $P\colon Y\to{\mathbb{C}}$ induces a homeomorphism $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ that restricts to a diffeomorphism on the set of non-exceptional orbits. Moreover, $q]\in{Y/\!/T}$ is exceptional exactly if $[q]=[1,0,0]$ and the degree $N$ of the defining polynomial of $Y$ is greater than $1$.
@@
Arguing as in the proof of Lemma 2.3, since $Y$ is tall the moment map $\Phi_{Y}$ is not proper. Hence, by [KT1, Corollary 6.3] the map $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is a homeomorphism. Moreover, by [KT1, Corollary 7.2] the restriction of $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ to the complement of the exceptional orbits is a diffeomorphism onto its image. By Lemma 2.9, if $q\in\Phi_{Y}^{-1}(0)$ is exceptional then $\overline{P}([q])=0$. Since $P([1,0,0])=0$, this implies that $[1,0,0]$ is the only possible exceptional orbit. Finally, by Lemma 2.9, $[1,0,0]$ is exceptional exactly if $N>1$. ∎
The next result allows us to understand the smooth structure on the reduced space Y//T{Y/\!/T} even when the homeomorphism $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is not a diffeomorphism.
@@
Let $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$ be a tall local model with defining monomial $P(z)=\prod_{i=0}^{h}z_{i}^{\xi_{i}}$. Assume that $\xi_{i}=0$ for all $i>k$. Let the group
@@
induces a diffeomorphism between ${\mathbb{C}}^{k+1}/\!/K$ and Y//T{Y/\!/T}.